Use the following cell phone airport data speeds​ (Mbps) from a particular network. Find P80.

0.2

0.2

0.3

0.3

0.3

0.4

0.4

0.4

0.5

0.6

0.7

0.8

0.8

0.9

0.9

0.9

1.1

1.1

1.3

1.3

1.4

2.1

2.3

2.5

2.7

2.7

2.9

3.2

3.8

4.5

5.4

5.8

7.5

8.3

8.5

8.8

9.5

9.6

9.6

9.7

10.2

10.4

12.9

13.5

14.3

15.6

15.6

15.6

15.7

25.8




1
Expert's answer
2022-06-16T09:43:59-0400

The data are in ascending order.

Compute the position of the pth percentile (index "i" ):


"i=\dfrac{p}{100}\times n=\dfrac{80}{100}\times50=40"

The index "i"  is an integer ⇒ the 80th percentile is the average of the values in the 40th and 41th positions (9.7 and 10.2 respectively)

Answer: the 80th percentile is (9.7+10.2) / 2 = 9.95